how many liters of a 45% salt solution should be mixed with 9 liters of a 35% salt solution to make a solution that is 36% salt? round to the nearest hundredth
Answer by ptaylor(2198)


Let x=number of liters of the 45% solution that is needed
Now we know that the pure salt in the 45% solution (0.45x) plus the amount of pure salt in the 35% solution (0.35*9) has to equal the amount of pure salt in the final mixture (0.36(9+x)). So, our equation to solve is:
0.45x+0.35*9=0.36(9+x) get rid of parens and simplify
0.45x+3.15=3.24+0.36x subtract 3.15 and also 0.36x from each side
0.45x-0.36x+3.15-3.15=3.24-3.15+0.36x-0.36x collect like terms
0.09x=0.09 divide each side by 0.09
x=1 liter-----------------------------amount of 45% solution needed
CK
0.45*1+0.35*9=0.36*10
0.45+3.15=3.60
3.60=3.60
Hope this helps----ptaylor